# Matched Filtering-Based Channel Estimation for AFDM Systems in Doubly Selective Channels

### v1

@@ II-C AFDM Demodulation
At the receiver side, the sample $r\left(n\right)$ is processed by DAFT to obtain $y\left(m\right)$ in the DAFT domain, i.e.,
$y\left(m\right)=\sum_{\bar{n}=0}^{N-1}{r}\left(\bar{n}\right)F_{c_{1},c_{2}}^{*}\left(\bar{n},m\right)+w\left(m\right),$ (10)
where $\left(\cdot\right)^{*}$ is the conjugate operation. The distribution of the noise $w\left(m\right)$ remains invariant. (10) can be rewritten as
$\boldsymbol{y}=\boldsymbol{Ar}=\boldsymbol{H}_{\mathrm{eff}}\boldsymbol{x}+\boldsymbol{w},$ (11)
where $\boldsymbol{y}=\left[y\left(0\right),\cdots,y\left(N-1\right)\right]^{\mathrm{T}}\in\mathbb{C}^{N\times 1}$, $\boldsymbol{H}_{\mathrm{eff}}=\sum_{i=1}^{P}{h_{i}}\boldsymbol{A\varGamma}_{\mathrm{CPP}_{i}}\boldsymbol{\varPi}^{l_{\tau_{i}}}\boldsymbol{\varDelta}^{k_{\nu_{i}}}\boldsymbol{A}^{\mathrm{H}}\in\mathbb{C}^{N\times N}$ is the effective channel matrix in the DAFT domain.
III Analysis of I/O, ICCI and SINR Loss
@@ III-A I/O Relation of AFDM
The $y\left(m\right)$ can be expressed as
$y(m)=\sum_{n=0}^{N-1}{x\left(n\right)h_{w}\left(m,n\right)}+w\left(n\right),$ (12)
$h_{w}\left(m,n\right)=\sum_{i=1}^{P}{h_{i}\mathcal{G}(m,n,k_{\nu_{i}},l_{\tau_{i}})},$ (13)
@@
and $\mathcal{F}\left(m,n,k_{\nu_{i}},l_{\tau_{i}}\right)$ is given by (64).
@@
Rewriting (12) into another matrix form, we have
$\boldsymbol{y}=\sum_{i=1}^{P}{h_{i}}\boldsymbol{H}_{i}\boldsymbol{x}+\boldsymbol{w},$ (16)
where $\boldsymbol{H}_{i}=\boldsymbol{A\varGamma}_{\mathrm{CPP}_{i}}\boldsymbol{\varPi}^{l_{\tau_{i}}}\boldsymbol{\varDelta}^{k_{\nu_{i}}}\boldsymbol{A}^{H}$.
III-B ICCI Analysis
@@
1) Integer Delay-Doppler (IDD) Channels: In this case, $l_{\tau_{i}}=\bar{l}_{\tau_{i}}$, $k_{\nu_{i}}=\bar{k}_{\nu_{i}}$. Then the I/O relation in (12) will be simplified to
$y(m)=\sum_{i=1}^{P}{h_{i}\alpha(m,n,k_{\nu_{i}},l_{\tau_{i}})x(n)}+w(n),$ (17)
@@
where $n=\left[m+2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}}\right]_{N}$, $\left[\cdot\right]_{N}$ is the mod-$N$ operation. For the $i$-th path, since only the transmitted signal $x\left(\left[m+2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}}\right]_{N}\right)$ constitutes $y\left(m\right)$, there is no ICCI for the IDD channels.
2) IDFD Channels: In this case, $l_{\tau_{i}}=\bar{l}_{\tau_{i}}$, $k_{\nu_{i}}=\bar{k}_{\nu_{i}}+\kappa_{\nu_{i}}$. Due to the existence of fractional Doppler $\kappa_{\boldsymbol{\nu}_{\boldsymbol{i}}}$, $\mathcal{F}(m,n,k_{\nu_{i}},l_{\tau_{i}})$ will not be zero for a given $m$ and any $n$. Some works have shown that $\left|\mathcal{F}(m,n,k_{\nu_{i}},l_{\tau_{i}})\right|$ has a peak at $n=\left[m+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}\right]_{N}$ and decreases significantly as $n$ moves away from $\left[m+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}\right]_{N}$. Therefore, we only need to consider the $12\xi+$ 1 principal values of $\mathcal{F}(m,n,k_{\nu_{i}},l_{\tau_{i}})$ around the peak of $\left[m+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}\right]_{N}$, i.e., $\left[m+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}-\xi\right]_{N}\leqslant n\leqslant\left[m+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}+\xi\right]_{N}$. Based on this good approximation, we can re-express $y\left(m\right)$ in (12) as
$y\left(m\right)\approx$ $\sum_{i=1}^{P}{\sum_{n=\left[m+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}-\xi\right]_{N}}^{\left[m+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}+\xi\right]_{N}}{x(n)h_{i}\mathcal{G}(m,n,k_{\nu_{i}},l_{\tau_{i}})}}$ (18)
$\approx$ $\sum_{i=1}^{P}{\sum_{q=-\xi}^{\xi}{x(n_{q})h_{i}\alpha(m,n_{q},k_{\nu_{i}},l_{\tau_{i}})}}$
×$\times\frac{e^{-j2\pi(q-\kappa_{\nu_{i}})}-1}{Ne^{-j\frac{2\pi}{N}(q-\kappa_{\nu_{i}})}-N},$
where $n_{q}=\left[m-q+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\boldsymbol{\nu}_{\boldsymbol{i}}}\right]_{N}$.
@@
3) FDFD Channels: In this case, $l_{\tau_{i}}=\bar{l}_{\tau_{i}}+\iota_{\tau_{i}}$, $k_{\nu_{i}}=\bar{k}_{\nu_{i}}+\kappa_{\nu_{i}}$. We rewrite $\mathcal{F}(m,n,k_{\nu_{i}},l_{\tau_{i}})$ as follows
$\mathcal{F}(m,n,k_{\nu_{i}},l_{\tau_{i}})=$ $\frac{1}{N}\frac{e^{-j2\pi(m-n+2Nc_{1}\left(\bar{l}_{\tau_{i}}+\iota_{\tau_{i}}\right)-\bar{k}_{\boldsymbol{\nu}_{\boldsymbol{i}}}-\kappa_{\boldsymbol{\nu}_{\boldsymbol{i}}})}-1}{e^{-j\frac{2\pi}{N}(m-n+2Nc_{1}\left(\bar{l}_{\tau_{i}}+\iota_{\tau_{i}}\right)-\bar{k}_{\boldsymbol{\nu}_{\boldsymbol{i}}}-\kappa_{\boldsymbol{\nu}_{\boldsymbol{i}}})}-1}$ (19)
$=$ $\frac{1}{N}\frac{e^{-j2\pi(m-n-\chi_{i})}-1}{e^{-j\frac{2\pi}{N}(m-n-\chi_{i})}-1}$
$=$ $\mathcal{J}(m,n,\chi_{i}),$
[…]
$y\left(m\right)\approx$ $\sum_{i=1}^{P}{\sum_{q=-\xi}^{\xi}{x\left(n_{q}\right)h_{i}\alpha\left(m,n_{q},k_{\nu_{i}},l_{\tau_{i}}\right)}}$ (20)
$\frac{e^{-j2\pi\left(q-\gamma_{i}\right)}-1}{Ne^{-j\frac{2\pi}{N}\left(q-\gamma_{i}\right)}-N}+w\left(m\right),$
@@
where $n_{q}=\left[m-q-\varUpsilon_{i}\right]_{N}$. From (18) and (20), it is evident that $y\left(m\right)$ is approximately a linear combination of $P\left(2\xi+1\right)$ transmitted signals. For the $12\xi+$ 1 transmitted signals in the $i$-th path, only $x\left(n_{q\left|q=0\right.}\right)$ is the main source of $y\left(m\right)$, and the other $2\xi$ signals can be regarded as interference. This interference is caused by the chirp-subcarriers near the $n_{q\left|q=0\right.}$-th chirp-subcarrier, so it is called ICCI.
III-C SINR
@@
$\hat{x}\left(n\right)=\left\{\boldsymbol{T}\right\}_{n,n}x\left(n\right)+\sum_{m=0,m\neq n}^{N}{\left\{\boldsymbol{T}\right\}_{n,m}}x\left(m\right)+\left\{\boldsymbol{Ww}\right\}_{n},$ (21)
where $n\in\left\{0,1,\cdots,N-1\right\}$, $\boldsymbol{T}=\boldsymbol{WH}_{\mathrm{eff}}$, {}$\left\{\cdot\right\}_{n,m}$ and {}$\left\{\cdot\right\}_{n}$ represent the elements at corresponding positions in the matrix or vector respectively. Therefore, the SINR on the $n$-th chirp-subcarrier is expressed as
$\mathrm{SINR}_{n}=\small{\frac{\left|\left\{\boldsymbol{T}\right\}_{n,n}\right|^{2}}{\sum_{m=0,m\neq n}^{N-1}{\left|\left\{\boldsymbol{T}\right\}_{n,m}\right|^{2}}+\sigma^{2}\left\|\left\{\boldsymbol{W}\right\}_{n,:}\right\|^{2}}}$ (22)
@@
III-D Effective SINR Loss
For the IDD channels, according to the I/O relationship of (17), the corresponding peak power of the $i$-th path is

### v3

@@ II-C AFDM Demodulation
At the receiver side, demodulation is performed to obtain $y\left[m\right]$ in the DAFT domain, i.e.,
$y\left[m\right]=\sum_{\bar{n}=0}^{N-1}{\left(r\left(t\right)g_{\mathrm{rx}}\left(t\right)\phi_{m}^{*}\left(t\right)\right)\left|{}_{t=\bar{n}T_{S}}\right.,}$ (12)
where $g_{\mathrm{rx}}\left(t\right)$ is the receive pulse-shaping filter. For simplicity, this paper adopts rectangular pulse shaping.
III Channel Analysis in The DAFT Domain
This section analyzes the I/O relationship, ICCI, SINR, and the effective SINR loss. It also presents two low-complexity methods for channel matrix construction.
@@ III-A I/O Relation of AFDM
The $y\left[m\right]$ can be expressed as
$y\left[m\right]=\sum_{n=0}^{N-1}{x\left[n\right]H_{\mathrm{eff}}\left[m,n\right]}+w\left[n\right],$ (13)
$H_{\mathrm{eff}}\left[m,n\right]=\sum_{i=1}^{P}{h_{i}\mathcal{G}(m,n,k_{\nu_{i}},l_{\tau_{i}})},$ (14)
@@
$\mathcal{F}\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)$ (17)
$=\frac{1}{N}\sum_{\bar{n}=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}}\left(\left(m-n+2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}}\right)\bar{n}-Nd_{\bar{n},n}\iota_{\tau_{i}}\right)}}$
[…]
@@
$\mathcal{G}(m,n,k_{\nu_{i}},l_{\tau_{i}})$ is the AFDM basis function used for subsequent CE. Rewriting (13) into another matrix form, we have
$\boldsymbol{y}=\boldsymbol{H}_{\mathrm{eff}}\boldsymbol{x}+\boldsymbol{w},$ (18)
where the effective channel matrix $\boldsymbol{H}_{\mathrm{eff}}$ is the matrix form of $H_{\mathrm{eff}}\left(m,n\right)$ in the DAFT domain. In addition, the permutation matrix can typically model the integer delay in the time-domain channel matrix $\boldsymbol{H}_{\mathrm{T}}$. However, this method fails for fractional delay. Since $\boldsymbol{H}_{\mathrm{eff}}=\boldsymbol{AH}_{\mathrm{T}}\boldsymbol{A}^{\mathrm{H}}$, $\boldsymbol{H}_{\mathrm{T}}$ is given by $\boldsymbol{H}_{\mathrm{T}}=\boldsymbol{A}^{\mathrm{H}}\boldsymbol{H}_{\mathrm{eff}}\boldsymbol{A}$.
III-B Analysis of ICCI and the Channel in the DAFT Domain
Since $\alpha\left(m,n,k_{\nu_{i}},l_{\tau_{i}}\right)$ is a phase factor, the characteristics of $\boldsymbol{H}_{\mathrm{eff}}$ are determined by $\mathcal{F}\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)$.
@@ III-B Analysis of ICCI and the Channel in the DAFT Domain
1) Integer-delay-integer-Doppler (IDID) Channels: In this case, $l_{\tau_{i}}=\bar{l}_{\tau_{i}}$, $k_{\nu_{i}}=\bar{k}_{\nu_{i}}$. Since $\mathcal{F}\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)$ simplifies to $\frac{1}{N}\delta\left(\left[m-n+2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}}\right]_{N}\right)$, (13) will be simplified to
$y\left[m\right]=\sum_{i=1}^{P}{h_{i}\alpha(m,n,k_{\nu_{i}},l_{\tau_{i}})x\left[n\right]}+w\left[n\right],$ (19)
@@
where $n=\left[m+2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}}\right]_{N}$. For the $i$-th path, since only the transmitted signal $x\left[\left[m+2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}}\right]_{N}\right]$ constitutes $y\left[m\right]$, there is no ICCI for the IDID channels.
@@
2) IDFD Channels: In this case, $l_{\tau_{i}}=\bar{l}_{\tau_{i}}$, $k_{\nu_{i}}=\bar{k}_{\nu_{i}}+\kappa_{\nu_{i}}$. Since $\iota_{\tau_{i}}=0$, we have
$\mathcal{F}\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)$ $=\frac{e^{\small{-j2\pi}\beta}-1}{Ne^{\small{\frac{-j2\pi}{N}}\beta}-1}$ (20)
$=\frac{1}{N}e^{-j(N-1)\pi\frac{\beta}{N}}\frac{\sin\left(\pi\beta\right)}{\sin\left(\frac{\pi\beta}{N}\right)},$
where $\beta=m-n+2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}}$. Due to the existence of fractional Doppler $\kappa_{\boldsymbol{\nu}_{\boldsymbol{i}}}$, $\mathcal{F}(m,n,k_{\nu_{i}},l_{\tau_{i}})$ will not be zero for a given $m$ and any $n$. Some works [6, 16] have shown that $\left|\mathcal{F}(m,n,k_{\nu_{i}},l_{\tau_{i}})\right|$ has a peak at $n=\left[m+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}\right]_{N}$ and decreases significantly as $n$ moves away from $\left[m+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}\right]_{N}$. Therefore, we only need to consider the $2\xi+1$ principal values of $\mathcal{F}(m,n,k_{\nu_{i}},l_{\tau_{i}})$ around the peak of $\left[m+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}\right]_{N}$, i.e., $\left[m+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}-\xi\right]_{N}\leqslant n\leqslant\left[m+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}+\xi\right]_{N}$. Based on this good approximation, we can re-express $y\left[m\right]$ in (13) as
$y\left[m\right]\approx$ $\sum_{i=1}^{P}{\sum_{n=\left[m+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}-\xi\right]_{N}}^{\left[m+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}+\xi\right]_{N}}{x\left[n\right]h_{i}\mathcal{G}(m,n,k_{\nu_{i}},l_{\tau_{i}})}}$ (21)
$\approx$ $\sum_{i=1}^{P}{\sum_{q=-\xi}^{\xi}{x\left[n_{q}\right]h_{i}\alpha(m,n_{q},k_{\nu_{i}},l_{\tau_{i}})}}$
$\times\frac{e^{-j2\pi(q-\kappa_{\nu_{i}})}-1}{Ne^{-j\frac{2\pi}{N}(q-\kappa_{\nu_{i}})}-N},$
@@
where $n_{q}=\left[m-q+2Nc_{1}l_{\tau_{i}}-\bar{k}_{{\nu}_{{i}}}\right]_{N}$. From (21), it is evident that $y\left[m\right]$ is approximately a linear combination of $P\left(2\xi+1\right)$ transmitted signals. For the $2\xi+1$ transmitted signals in the $i$-th path, only $x\left[n_{q\left|q=0\right.}\right]$ is the main source of $y\left[m\right]$, and the other $2\xi$ signals can be regarded as interference. This interference is caused by the chirp-subcarriers near the $n_{q\left|q=0\right.}$-th chirp-subcarrier, so it is called ICCI.
3) FDFD Channels: In this case, $l_{\tau_{i}}=\bar{l}_{\tau_{i}}+\iota_{\tau_{i}}$, $k_{\nu_{i}}=\bar{k}_{\nu_{i}}+\kappa_{\nu_{i}}$. Since $\iota_{\tau_{i}}\neq 0$, we have
$\mathcal{F}\left(m,n,k_{\nu_{i}},l_{\tau_{i}}\right)=\frac{1}{N}\sum_{\bar{n}=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}}\eta_{m,n}^{\left(i\right)}\bar{n}}e^{j2\pi\left(d_{\bar{n},n}-2\bar{n}c_{1}\right)\iota_{\tau_{i}}},}$ (22)
where $\eta_{m,n}^{\left(i\right)}=m-n-\chi_{i}-\kappa_{\nu_{i}}$, $\chi_{i}=-2Nc_{1}\bar{l}_{\tau_{i}}+\bar{k}_{\nu_{i}}$ is the equivalent shift of the $i$-th path. Since $\left(d_{\bar{n},n}-2\bar{n}c_{1}\right)\iota_{\tau_{i}}$ is small, it does not affect the peak position. For the $i$-th path, the condition for the occurrence of the main peak of $\left|\mathcal{F}\left(m,n,k_{\nu_{i}},l_{\tau_{i}}\right)\right|$ is
$m=m_{\mathrm{m-peak}}^{\left(i\right)}=\left[n+\chi_{i}\right]_{N}.$ (23)
In addition, $\left|\mathcal{F}\left(m,n,k_{\nu_{i}},l_{\tau_{i}}\right)\right|$ displays some high-level local peaks. The positions of these local peaks are
$m_{\mathrm{l}-\mathrm{peak}}^{\left(i\right)}=m_{\mathrm{m}-\mathrm{peak}}^{\left(i\right)}\pm 2gNc_{1},$ (24)
where $g$ is a small positive integer. In addition, as $m_{\mathrm{l}-\mathrm{peak}}^{\left(i\right)}$ moves away from $m_{\mathrm{m}-\mathrm{peak}}^{\left(i\right)}$, its local peaks decrease in magnitude. Therefore, for FDFD channels, these local peaks will lead to more severe ICCI. Simultaneously, local peaks in all paths further exacerbate the inter-path interference.
Based on the above analysis, since $\left|\mathcal{F}\left(m,n,k_{\nu_{i}},l_{\tau_{i}}\right)\right|$ contains non-negligible local peaks, the I/O relationship cannot be approximated and can only be expressed as (13). In addition, N1N-1 transmitted signals will generate ICCI for $y\left[m\right].$
III-C Low-Complexity Channel Matrix Construction
The rapid and low-complexity construction of channel matrices enables efficient system simulations, reduces computational overhead, and facilitates real-time signal processing, which is particularly important for high-dimensional or large-scale systems.
The complexity of constructing $\boldsymbol{H}_{\mathrm{eff}}$ mainly comes from the construction of the $\mathcal{F}\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)$. For both IDID and IDFD channels, $\mathcal{F}\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)$ can be converted into the simple form of the summation-free version in (17). However, for FDFD channels, due to the presence of $d_{\bar{n},n}\iota_{\tau_{i}}$, $\mathcal{F}\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)$ cannot be further simplified. Using the element-wise summation (ES) method to construct the channel matrix leads to high complexity, especially for large $N$.
1) The FFT-based Method:
We rewrite $\mathcal{F}\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)$ as
$\mathcal{F}\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)$ $=\sum_{\bar{n}=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}}\left(\left(2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}}-n\right)\bar{n}-Nd_{\bar{n},n}\iota_{\tau_{i}}\right)}}$ (25)
$\hskip 30.0pt\times e^{\small{\frac{-j2\pi}{N}}m\bar{n}}$
$=\sum_{\bar{n}=0}^{N-1}{u\left(\bar{n},n\right)e^{\small{\frac{-j2\pi}{N}}m\bar{n}}},$
where $u\left(\bar{n},n\right)=e^{\small{\frac{-j2\pi}{N}}\left(\left(2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}}-n\right)\bar{n}-Nd_{\bar{n},n}\iota_{\tau_{i}}\right)}$. This is exactly the DFT Transform. Utilizing the FFT enables the rapid construction of $\mathcal{F}\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)$.
2) The Segment-wise Summation (SS) Method:
Considering the special structure of $d_{\bar{n},n}$, $\left|\mathcal{F}(m,n,k_{\nu_{i}},l_{\tau_{i}})\right|$ can be simplified into a relatively simple segment-wise summation form. $d_{\bar{n},n}$ exhibits a staircase structure. For a given fixed $n$, the staircase has $R_{n}$ levels. In addition, let the starting index, width, and value of each level are $\bar{n}_{r,n}$, $L_{r,n}$, and $d_{r,n}$, respectively. Therefore, (22) can be rewritten as
$\mathcal{F}\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)$ $=\frac{1}{N}\sum_{r=0}^{R_{n}-1}{e^{j2\pi d_{r,n}\iota_{\tau_{i}}}\sum_{\bar{n}=\bar{n}_{r,n}}^{\bar{n}_{r,n}+L_{r,n}-1}{e^{\small{\frac{-j2\pi}{N}}\beta\bar{n}}}}$ (26)
$=\frac{1}{N\sin\left(\frac{\pi\beta}{N}\right)}\sum_{r=0}^{R_{n}-1}{\sin\left(\frac{\pi}{N}\beta L_{r,n}\right)e^{j\varphi_{r,n}}},$
where $\varphi_{r,n}=2\pi d_{r,n}\iota_{\tau_{i}}-\small{\frac{2\pi\beta}{N}\bar{n}_{r,n}^{\mathrm{c}}}$, $\bar{n}_{r,n}^{\mathrm{c}}=\bar{n}_{r,n}+\frac{L_{r,n}-1}{2}$ is the central index of the $r$-th segment, $\beta=m-n+2Nc_{1}l_{\tau_{i}}-k_{\nu_{i}}$. In addition, the mean $R$ of $R_{n}$ is generally $R\approx 2Nc_{1}$.
The ES method directly calculates each matrix element through explicit summation over all $\bar{n}$, resulting in a cubic complexity of $\mathcal{O}\left(N^{3}\right)$. The FFT-based method exploits the inherent DFT structure of the summation over $\bar{n}$ in $\mathcal{F}\left(m,n,l_{\tau_{i}},k_{\nu_{i}}\right)$. By applying FFT to accelerate the computation, the overall complexity is reduced to $\mathcal{O}\left(N^{2}\log N\right)$. By leveraging the staircase structure of $d_{\bar{n},n}$, the SS method decomposes the summation over $n$ into a series of closed expressions from each segment. Thus, the SS method effectively reduces the summation size from $N$ to $R$, yielding complexity $\mathcal{O}\left(N^{2}R\right)$, far below the complexity of ES method.
III-D SINR
@@
$\hat{x}\left[n\right]=T\left[n,n\right]x\left[n\right]+\sum_{m=0,m\neq n}^{N}{T\left[n,m\right]x\left[m\right]}+\tilde{w}\left[n\right],$ (27)
where $n\in\left\{0,1,\cdots,N-1\right\}$, $\boldsymbol{T}=\boldsymbol{WH}_{\mathrm{eff}}$. Therefore, the SINR on the $n$-th chirp-subcarrier is expressed as
$\mathrm{SINR}_{n}=\small{\frac{\left|T\left[n,n\right]\right|^{2}}{\sum_{m=0,m\neq n}^{N-1}{\left|T\left[n,m\right]\right|^{2}}+\sigma^{2}\left\|\left\{\boldsymbol{W}\right\}_{n,:}\right\|^{2}}},$ (28)
@@
III-E Effective SINR Loss
For the IDID channels, according to the I/O relationship of (19), the corresponding peak power of the $i$-th path is